The factory owner claimed that their bottled fruit juice has the capacity of less than an average of 280ml. To test the claim, a group of consumers gets a sample of 80 bottles of the fruit juice, calculates the capacity, and then finds the mean capacity to be 265ml. The standard deviation is 8ml. Use a= 0.05 level of significance to test the claim.

1
Expert's answer
2022-06-02T17:19:17-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge280"

"H_1:\mu<280"


3. This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


4. Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z<-1.6449\}."


5. The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{265-280}{8/\sqrt{80}}=-16.7705"


6. Since it is observed that "z=-16.7705<-1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed is "p=P(Z<-16.7705)=0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 280, at the "\alpha = 0.05" significance level.